### v1

# Molecular Fingerprints Are Strong Models for Peptide Function Prediction

Abstract
We study the effectiveness of molecular fingerprints for peptide property prediction and demonstrate that domain-specific feature extraction from molecular graphs can outperform complex and computationally expensive models such as GNNs, pretrained sequence-based transformers and multimodal ensembles, even without hyperparameter tuning. To this end, we perform a thorough evaluation on 126 datasets, achieving state-of-the-art results on LRGB and 5 other peptide function prediction benchmarks. We show that models based on count variants of ECFP, Topological Torsion, and RDKit molecular fingerprints and LightGBM as classification head are remarkably robust. The strong performance of molecular fingerprints, which are intrinsically very short-range feature encoders, challenges the presumed importance of long-range interactions in peptides. Our conclusion is that the use of molecular fingerprints for larger molecules, such as peptides, can be a computationally feasible, low-parameter, and versatile alternative to sophisticated deep learning models.
Machine Learning, ICML
1 Introduction
Peptides are small proteins, typically with amino acid chains of 3333-$5050$ residues and hundreds of atoms (Singh et al., 2015). They perform a variety of biological functions and have attracted increasing attention from researchers, mainly in the context of peptide-based drugs. Their useful therapeutic properties include antiviral, antitoxin, and combating cancer cells (Goles et al., 2024). They are also exceptionally effective in treating obesity and diabetes (Wang et al., 2022). In particular, antimicrobial peptides (AMPs) offer a potential solution to the escalating antimicrobial resistance crisis by offering an alternative to traditional antibiotics (Szymczak et al., 2023; Santos-Júnior et al., 2024).
Peptide function prediction problems are not only important in practice, but also manifest specific challenges in building machine learning pipelines, such as negative class sampling (Sidorczuk et al., 2022) or homology bias (Fernández-Díaz et al., 2024). This motivated the creation of multiple benchmark datasets that promote fair comparison with predefined splits, varied targets, and established evaluation protocols.
Peptides can be represented in many different modalities, with the selected representation determining the applicable machine learning models. When stable spatial folding structure is available, structural encodings can be used (Spänig et al., 2021), but this is relatively rare for peptides. In bioinformatics, typical approaches rely on complex feature engineering in amino acid sequences (Sidorczuk et al., 2022), with newer methods leveraging pretrained Protein Language Models (PLMs) (Elnaggar et al., 2022). Alternatively, peptides can be represented at the atom level by utilizing their topological molecular graphs (without modeling the spatial structure). This approach has been particularly popular in the machine learning literature due to the rapid development of Graph Neural Networks (GNNs) and Graph Transformers (GTs). Long Range Graph Benchmark (LRGB) (Dwivedi et al., 2022) in particular proposed the peptide function prediction as a graph-level, long-range interaction problem. It has been extensively used to benchmark the ability of novel GNN architectures to effectively capture long range dependencies between graph nodes.
In chemoinformatics, molecular fingerprints (Todeschini & Consonni, 2009) are one of the most popular models for feature extraction from small molecules, widely applied to similarity search, chemical space exploration, and molecular property prediction (Muegge & Mukherjee, 2016). They have also shown remarkable performance compared to GNNs (Jiang et al., 2021; Baptista et al., 2022; Xia et al., 2023), indicating the importance of domain-specific encoding algorithms. However, the adoption of this approach to predict peptide functions remains very limited, with just a few hybrid models that incorporate binary fingerprints as one component of their approach (Kang et al., 2024; Tan et al., 2024; Liang et al., 2024). Recent studies (Jiang et al., 2021; Tönshoff et al., 2024; Adamczyk & Czech, 2024) suggest that domain-specific features, paired with a robust nonlinear classifier, can surpass complex models in performance on both small and large molecules, achieving impressive results also on the LRGB benchmark.
In this study, we revisit the effectiveness of count-based variants of hashed molecular fingerprints for predicting peptide functions. These models, by design, are highly localized, counting occurrences of small, short-range subgraphs. When combined with a LightGBM classifier, fingerprint representations achieve state-of-the-art performance on the LRGB benchmark as well as five additional peptide function prediction benchmarks, spanning a total of 126 datasets.
The key contributions of this work are as follows. First, we show that count-based variants of hashed molecular fingerprints, coupled with the LightGBM classifier, obtain SOTA results in varied peptide function prediction tasks, with other advantages, such as simplicity and computational efficiency. We benchmark this method on the largest (to our knowledge) collection of 6 benchmarks and 126 datasets. Secondly, we show that these inherently very short-range models outperform long-range GNNs, challenging the presumed importance of long-range interactions in peptides and their viability for benchmarking long-range graph models.
2 Related works
The wide variety of prediction models for peptide functions arises from differences in input representations (e.g. sequence-based vs. graph-based, structural encodings, or feature extraction vs. learned representations) and reflects the influence of distinct research communities and the diversity of underlying datasets.
Considering granular, atom-level peptide representation in a form of molecular graph, Graph Neural Networks (GNNs) provide an automated method for building prediction models capable of incorporating not only structural information but also atom and bond attributes. The field of graph representation learning has advanced significantly since the introduction of foundational models such as Graph Convolutional Network (GCN) (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017) or Graph Isomorphism Network (GIN) (Xu et al., 2019), paving the way for more expressive models such as GINE (Hu et al., 2020) or GatedGCN (Dwivedi et al., 2023). However, these GNNs, limited by their reliance on iterative one-hop message passing, struggle to effectively capture information from distant nodes, which can be crucial for classification of larger molecules. In contrast, Graph Transformers (GT) overcome this limitation by allowing direct attention between all nodes, yet their lack of inherent graph inductive biases forces them to depend on arbitrary positional encodings, undermining their ability to fully exploit graph structure. The tradeoff between those two factors led to the development of modern hybrid GNN architectures, which display much better accuracy on molecular benchmarks. GraphGPS (Rampášek et al., 2022) combines local message passing with global attention and positional encodings, helping the attention mechanism understand the relative positions of the nodes. Enhanced using the Hierarchical Distance Structural Encoding (HDSE) mechanism (Luo et al., 2024), GraphGPS achieves one of the best results in the Peptide-func dataset from LRGB. In (Gutteridge et al., 2023), the Dynamically Rewired Message Passing (DRew) model was introduced. It is an adaptable framework, for any message-passing GNN, incorporating layer-dependent rewiring, controlled message passing delays, and skipping connections. On the other hand, the graph inductive bias transformer (GRIT) (Ma et al., 2023) does not require message passing, but still captures local and global graph properties using degree information injection and learned relative positional encodings. Subsequently, other approaches have been proposed to capture long-range dependencies. Graph ViT (He et al., 2023) is a unique model for graph representation learning that integrates the strengths of GNN-based encoders for local feature extraction with Vision Transformers to capture global context effectively. Finally, S2GNNs (Spatio-Spectral Graph Neural Networks) (Geisler et al., 2024) are a novel type of graph neural network that combines spatial message passing with spectral graph filters and have recently been shown to outperform state-of-the-art graph transformers in peptide function prediction tasks.
Due to the limited data availability, many peptide prediction tasks involve only around a thousand samples or fewer. Furthermore, simple features like amino acid compositions are known to be strong predictors of many properties. As a result, the use of deep learning remains limited in bioinformatics tasks on peptides. Instead, numerous feature extraction methods, often quite sophisticated, have been developed. They often utilize multiple modalities, extracting physicochemical properties of peptide molecules (e.g. amphiphilicity, hydrophobicity), as well as compositional features from amino acid sequences. Such methods include MLAMP (Lin & Xu, 2016), ampir (Fingerhut et al., 2020), and MACREL (Santos-Júnior et al., 2020) or amPEPpy (Lawrence et al., 2021). Combined with traditional classifiers such as SVM and Random Forest, they have shown good performance, even compared to deep learning methods like AMPScannerV2 (Veltri et al., 2018).
Novel deep learning models on amino acid sequences are based on pretrained transformers. Those Protein Language Models (PLMs) are trained on large-scale datasets with proteins typically much larger than peptides. They are heavily inspired by natural language processing (NLP) and include, e.g. ProtBERT and Prot-T5-XL (Elnaggar et al., 2022). The authors of the ESM architecture (Lin et al., 2023) showed that they exhibit a scaling behavior similar to large language models (LLMs) in NLP. ESM models have also shown particularly strong performance in independent benchmarks compared to other PLMs (Fernández-Díaz et al., 2024).
Since antimicrobial peptides (AMPs) are particularly important in drug design, there are many specialized PLMs for predicting this property, including BERT-Protein (Zhang et al., 2021), LMPred (Dee, 2022), cAMPs_pred (Ma et al., 2022), and AMP-BERT (Lee et al., 2023). They can be used out-of-the-box with pretrained AMP classification heads.
3 Methods
Here, we describe the molecular fingerprints that form our proposed atom-level peptide encoding. We also describe the design of Long Range Graph Benchmark and their approach to testing the presence of long-range dependencies.
3.1 Molecular fingerprints
Molecular fingerprints are feature extraction algorithms for molecules. The vast majority of them are based on extracting small subgraphs and detecting their presence or counting occurrences (binary and count variants, respectively). This turns the problem of molecular graph classification into tabular classification, typically combined with tree-based ensemble classifiers such as Random Forest (Breiman, 2001) or LightGBM (Ke et al., 2017).
We focus on hashed fingerprints (Todeschini & Consonni, 2009), exemplified by ECFP (Rogers & Hahn, 2010) or Topological Torsion (Nilakantan et al., 1987). They are more flexible than structural fingerprints like MACCS or PubChem, which use predefined subgraphs like functional groups or rings of given size, and are not designed for peptides. Hashed fingerprints use a more flexible approach, defining a general “shape” of extracted subgraphs, for example, circular atom neighborhoods in ECFP (Rogers & Hahn, 2010), short paths in Topological Torsion (TT) (Nilakantan et al., 1987), or all small subgraphs in RDKit fingerprint (rdk, ). Extracting subgraph features in this manner offers greater flexibility, as it does not, for instance, rely on specific atom types. Subgraphs are defined by their structure, including the specific atoms and bonds they contain (e.g. characterized by atomic numbers and bond types), which are then converted into a unique integer identifier. To obtain a constant-length feature vector, they are hashed into the output vector, using the modulo function in the simplest case. The binary variant simply indicates whether a given substructure appears in a molecule at all, while the count-based variant tallies the occurrences, also accounting for potential hashing collisions.
ECFP typically uses circular subgraphs of radius 2 (see Figure 1). It starts with the atom itself, creating a numerical identifier based on atomic number, number of heavy neighbors, valence, and a few other simple features (see (Rogers & Hahn, 2010) for a complete list). Then, it iteratively increases the radius, incorporating neighbors by combining their identifiers, up to a given radius. This results in the final subgraph identifier for each atom. Such a procedure, known as the Morgan algorithm, closely resembles the operation of message passing in GNNs, as both are rooted in the Weisfeiler-Lehman isomorphism test. As such, ECFP with radius 2 is roughly equivalent to a shallow 2-layer GNN. Topological Torsion (Nilakantan et al., 1987), designed to model short-range molecular interactions, encodes linear paths of length 4, combining the identifiers of these consecutive atoms. RDKit fingerprint (rdk, ) uses all subgraphs of size up to 7 bonds, which can be nonlinear and include cycles (e.g. rings). The major difference between TT and RDKit is that the former uses only linear paths, while the RDKit fingerprint also encodes small cyclic structures (e.g. benzene rings).
We focus on those three fingerprints, as they are strictly local, very short-range descriptors. Therefore, a good performance of models based on their features can be attributed only to the importance of the short-range relations, rather than to long-range dependencies, since they are inherently incapable of capturing them. This locality makes them notably fast to compute (Adamczyk & Ludynia, 2024) and scalable to larger molecules like peptides, e.g. ECFP scales linearly with the number of atoms. Because peptides are relatively linear structures, path-based TT and RDKit fingerprints also scale well. Those three fingerprints have also shown remarkable performance in ligand-based virtual screening (Riniker & Landrum, 2013).
Figure 1: Example subgraph extraction in consecutive iterations of ECFP algorithm. Reprinted with permission from (Rogers & Hahn, 2010). Copyright 2010 American Chemical Society.
3.2 Long-range interactions (LRI)
Long Range Graph Benchmark (LRGB) (Dwivedi et al., 2022) proposed 3 conditions for a dataset to potentially require long-range interactions (LRI). Firstly, a necessary condition is that the graphs have to be sufficiently large, in particular larger than typical GNNs receptive field. Secondly, the nature of the task itself has to indicate that long-range dependencies between nodes are crucial to properly predict properties, with local interactions not being enough. Lastly, the experimental results should support this, i.e. models that incorporate global knowledge, like GNNs with topological encodings or graph transformers, have to outperform local message-passing ones. For graph classification and regression, peptide property prediction was proposed as a group of tasks that meet these conditions.
We show that this is not necessarily the case, particularly in relation to the third assumption. Peptides at the atomic level are certainly large and “long” (large diameter, low atom degrees) enough graphs to suggest a possible LRI. However, the biological motivation for LRI in peptides is somewhat arguable. Proteins have long-range dependencies because of their large structures, which enable complex folding through distant amino acid interactions. In contrast, the smaller size of the peptides naturally limits such possibilities.
The third point, i.e. the superior performance of models with global knowledge, is the easiest to validate quantitatively. Since the argument is based on experimental results, it suffices to show that inherently local short-range models outperform long-range ones. Molecular fingerprints exactly fulfill these requirements. By obtaining SOTA results, we show that they provide strong evidence against the third claim.
4 Experiments and results
We train proposed fingerprint-based models on a variety of benchmarks that cover a wide range of peptide properties, data sampling methods, quality metrics, and train-test splits. We adhere closely to the procedure outlined for each benchmark in its original publication and report the corresponding quality metrics. For benchmarks that provide multiple metrics, we present AUROC and Matthews Correlation Coefficient (MCC) here due to space constraints, with the remaining metrics detailed in Appendix A. The best result for each table is highlighted in bold.
Our proposed fingerprint-based models always use their count variants and are implemented with the scikit-fingerprints (Adamczyk & Ludynia, 2024) library. We use default values for all hyperparameters: ECFP radius 2 (diameter 4, also known as ECFP4), Topological Torsion path length 4, and RDKit fingerprint path length 7. They use small subgraphs as features, capturing only very short-range dependencies. We did not see any significant improvement when tuning those hyperparameters (except for PeptideReactor, see Section 4.3).
We use LightGBM (Ke et al., 2017) with 500 trees as the classification model, due to its robust performance. Since most datasets are heavily imbalanced, we use class weight balancing inversely proportional to the frequency of the positive class. We keep default values for all other hyperparameters because we noticed negligible improvements when tuning them. Only setting the number of trees higher than the default 100 was relevant for performance, which is consistent with the literature (Probst et al., 2019).
4.1 Long Range Graph Benchmark (LRGB)
To begin, we report the results on the LRGB (Dwivedi et al., 2022) benchmark, on both Peptides-func (classification) and Peptides-struct (regression), in Table 1. Those are multioutput datasets, and we report the metric values averaged from all tasks. We include original Graph Transformer and SAN from LRGB (Dwivedi et al., 2022), as well as a variety of recently published models, including MOLTOP baseline (Adamczyk & Czech, 2024), well-tuned classic GNNs (GCN, GatedGCN, GINE) (Tönshoff et al., 2024), and various long-range models such as GRIT (Ma et al., 2023), HDSE (Luo et al., 2024) and S2GCN (Geisler et al., 2024).
Since both datasets are multitask and LightGBM only supports single outputs, we train a separate model per task and combine their predictions. For Peptides-struct, we optimize MAE as a loss function. LRGB results for GNNs are reported for 10 different random seeds. However, LightGBM with default settings is a deterministic classifier, so we don’t report its standard deviation (since it would be zero). In Appendix B, we include results for the Random Forest and Extremely Randomized Trees classifiers, which depend on random seed. However, their standard deviations are very low (<0.1absent0.1<0.1< 0.1 AUPRC).
[…]
ECFP4 fingerprint is roughly equivalent to a 2-layer message-passing GNN. However, it outperforms both local GNNs such as GCN and long-range models like S2GCN and HDSE. We postulate that the reason for this difference is three-fold. First, fingerprints are based on counting discrete substructures, rather than learning continuous graph representations, resulting in a stronger inductive bias. This reduces dependence on training data, as GNNs must learn everything from scratch. Secondly, peptides are built from repeatable fragments, as amino acids are structurally similar and often occur multiple times. Therefore, GNNs may primarily need to recognize that certain small subgraphs, such as the amine group -NH2, frequently occur in peptides. This phenomenon is not indicative of long-range interaction or dependency, but rather reflects the straightforward observation that common subgraphs appear repeatedly across different regions of the input. Third, count-based fingerprints capture the influence of the structure size and frequency of particular substructures. Regression targets in Peptides-struct like inertia mass and length are strongly related to peptide size, and classification labels (e.g. antimicrobial) depend on binding properties of peptides, which also differ with available surface area and thus size. This means that they may not be related to any long-range interactions, but rather local structures and overall peptide size, which can be well approximated by counting subgraphs.
Therefore, our experimental results lead us to believe that Peptides-func and Peptides-struct datasets do not exhibit long-range interactions or require learning long-range dependencies. This has been overlooked in previous works because in (Dwivedi et al., 2022) the authors used only GNNs, and the re-assessment of (Tönshoff et al., 2024) included only quite deep models (6-10 layers) with global topological encodings. This highlights the importance of comparing against established, domain-specific models for fair evaluation. To highlight the impact of using count-based fingerprints, we also compare them with classifiers trained on binary variants, with differences summarized in Table 2. We observe significant performance gains when using count features instead of binary ones. In light of this finding, it is concerning that all the hybrid models in the bioinformatics literature we reviewed rely exclusively on binary fingerprints, especially since count variants incur the same computational cost and result in significantly stronger models.
Table 1: Results on LRGB benchmark.
Model
Peptides-func
AUPRC $\uparrow$
Peptides-struct
MAE $\downarrow$
[…]
RDKit 73.11 0.2459
$T$ 73.18 0.2438
ECFP 74.60 0.2432
We also note the efficiency of our approach. Treating a LightGBM tree node as one parameter, the models for all fingerprints use approximately 300k parameters, in total, well below the budget of 500k set by LRGB. In terms of wall time, computing ECFP fingerprints and training LightGBM takes 19 seconds on a 12-core Intel Core i7-12700KF CPU on Peptides-func. In comparison, training graph transformer or SAN takes up to 60 hours on NVidia A100 GPU (Dwivedi et al., 2022). We provide additional timings in Appendix C.
Table 2: Differences between binary and count fingerprint models.
Dataset Variant ECFP $T$ RDKit
Peptides-func binary 70.57 66.18 63.88
Peptides-func count 74.60 73.18 73.11
AUPRC gain $\uparrow$ +4.03 +7.00 +9.23
Peptides-struct binary 0.3049 0.3298 0.3331
Peptides-struct count 0.2432 0.2438 0.2459
MAE gain $\downarrow$ -0.0617 -0.0860 -0.0872
4.2 Antimicrobial peptides (AMPs) benchmarks
In the next experiment, we investigate whether the results obtained on the LRGB generalize to other benchmarks in this domain. Such a generalization would suggest that predicting peptide function is not inherently reliant on long-range interactions.
We present results on antimicrobial peptide (AMP) prediction benchmarks. Although the target remains the same (to distinguish peptides with antimicrobial properties), the benchmarks differ in the methodology of selecting positive and negative data, structural diversity, and the way of performing the train-test split (e.g. incorporating sequence homology or clustering). We select 3 benchmarks, covering varied approaches and procedures. Note that many of the datasets used in this section call the data “peptides”, also if those proteins have on average, e.g. 70, 100, or even 200 amino acids. Therefore, to achieve strong performance, fingerprint-based models must be both scalable (owing to their operation at the atomic level) and capable of accurately predicting the properties of larger molecules.
First, we compare our approach with various BERT-like models pretrained for AMP prediction, as evaluated in (Gao et al., 2024). They include AMP peptides from 6 classic datasets, with negative samples taken from UniProt, and additionally filtering the dataset using CD-HIT (Fu et al., 2012) at similarity threshold 40% to remove homology bias. The authors infer only with BERT models, since they have already been pretrained and fine-tuned for AMP prediction. They also note that parts of those datasets may even be in the pretraining datasets, to the advantage of said models on this benchmark. To simulate such pretraining with fingerprints and LightGBM, we use the leave-one-out-dataset strategy. For example, when testing on ADAPTABLE dataset, we merge all other datasets into one training set, and apply CD-HIT-2D (Fu et al., 2012) (threshold 40%) to perform homology filtering, i.e. remove peptides too similar to the ones in the dataset currently used as test set.
[…]
Table 3: Results on AMPs benchmark from (Gao et al., 2024).
Model ADAPTABLE APD CAMP dbAMP DRAMP YADAMP Average F1
AMP-BERT 81.7 78.3 82.3 62.7 64.0 82.6 75.3
BERT-Protein 79.4 90.1 80.6 82.9 58.8 58.8 75.1
cAMPs_pred 61.3 74.1 72.8 52.4 39.2 85.0 64.1
LM_pred 55.9 71.7 64.5 60.7 47.6 82.1 63.8
LM_pred (BFD) 73.9 89.6 83.6 68.7 64.8 84.4 70.3
RDKit 82.0 93.0 88.4 91.4 76.4 92.0 87.2
$T$ 80.6 94.3 88.8 90.8 76.9 95.8 87.8
ECFP 80.8 93.0 88.2 91.5 76.3 95.8 87.6
Still, those 6 datasets are relatively small. The authors of (Xu et al., 2021) created a unified AMP benchmarking dataset based on 9 commonly used smaller ones (e.g. CAMP and DRAMP), merged and deduplicated using homology-based CD-HIT clustering (Fu et al., 2012). This ensures a high-quality and hard train-test split by avoiding the presence of too similar structures. They benchmarked traditional feature engineering methods, combined with different classifiers (e.g. Random Forest, SVM, kNN).
The comparison is presented in Table 4. Fingerprints achieve SOTA performance, indicating that they perform well at different data sizes. The best previous method, AMPfun (Chung et al., 2019), is a relatively complex pipeline, using binary profiling of amino acid positions, amino acid composition characteristics, and physicochemical descriptors, combined with feature selection. Such approaches can be relatively fragile, as demonstrated by the superior performance of much more straightforward ECFP in both AUROC and MCC metrics.
Table 4: Results on AMPs benchmark from (Xu et al., 2021).
Method AUROC MCC
AMPscannerV2 58.5 0.137
iAMP-2L 59.2 0.261
ADAM-SVM 61.2 0.264
ampir 61.9 0.156
MLAMP 62.9 0.194
ADAM-HMM 68.4 0.39
AMPlify 69.7 0.381
AMPEP 72.7 0.425
AMPfun 73.5 0.414
RDKit 74.5 0.404
$T$ 73.1 0.385
ECFP 75.0 0.429
As noted in (Sidorczuk et al., 2022), the negative class sampling has a considerable effect on the model performance. Their AMPBenchmark measures the sensitivity of various AMP models to the choice of data by retraining the models on a large collection of datasets and varying negative class sampling methods. Models vary from sequence-based feature engineering (e.g. AmpGram (Burdukiewicz et al., 2020)), CNN-based deep learning (e.g. Deep-AmPEP30 (Yan et al., 2020)), to multimodal ensembles utilizing physicochemical and amino acid features (e.g. ampir (Fingerhut et al., 2020)).
AMPBenchmark tests the robustness of the models by retraining them on a collection of datasets. The same positive class peptides are combined with 11 different negative class sampling approaches, each repeated 5 times with different data from the UniProt database. This results in 55 datasets, differing in amino acid sequence lengths, sequence similarities, and class balance. Each uses 80-20% train-test split. The results obtained on the AMPBenchmark are summarized in Table 5, including the mean and standard deviation of retraining on different datasets and negative sampling sets. Again, all fingerprint-based models achieve the best results, and with standard deviations considerably lower than other methods. This indicates that feature engineering based on a peptide topological graph may result in more robust and reliable models compared to complex ensembles. Although popular in the bioinformatics literature, they are inherently prone to human bias in the choice of descriptors.
Table 5: Results on AMPs benchmark from (Sidorczuk et al., 2022).
Model AUROC
[…]
4.3 General peptide benchmarks
In addition to antimicrobial properties, there are numerous intriguing peptide attributes to predict, such as antioxidant activity, anti-MRSA effects, toxicity, and others. However, a common challenge in this domain is the limited size of the dataset, often containing fewer than 1,000 positive samples per dataset. To evaluate the performance of fingerprint-based models under these constraints, we selected two comprehensive benchmarks that encompass a total of 68 datasets that cover a diverse range of peptide properties.
AutoPeptideML (Fernández-Díaz et al., 2024) consists of 18 datasets. Their positive peptides have been taken from the dataset available in the literature, but the negative samples and the train-test split utilize a novel NegSearch+HP approach to ensure a particularly challenging setup. Negative peptides are taken from APML-Peptipedia, filtered to avoid false negatives, and sampled to have similar sequence length distribution to positive ones. The train-test split uses a homology clustering algorithm to create an out-of-distribution test set, somewhat similar to the scaffold split, popular in molecular property prediction. AutoPeptideML provides results on state-of-the-art PLMs, including LLMs with 3B weights such as Prot-T5-XL (Elnaggar et al., 2022).
The results are summarized in Table 6. Although fingerprint-based approaches do not achieve the highest overall performance, they are close in terms of MCC. Moreover, they are significantly more parameter-efficient, achieving comparable performance with only about 22k parameters, in contrast to PLMs, which can have up to 3 billion weights. In particular, the ECFP-based classifier has better AUROC and MCC than three variants of ESM2 (Lin et al., 2023), which are known to be a particularly powerful protein model, and MCC almost identical to ESM2-650M. The reason for this may be that peptides are very short amino acid sequences, compared to most proteins, which were the bulk of the training and testing data for PLMs. Fingerprints operate on larger molecular graphs, which can also be seen as “higher-resolution" peptide representation than sequences.
Table 6: Results on AutoPeptideML benchmark (Fernández-Díaz et al., 2024).
Model # params AUROC MCC
ProtBERT 420M 75.9 0.375
ESM2-150M 150M 77.7 0.402
Prost-T5 3B 77.1 0.409
ESM2-8M 8M 77.5 0.418
ESM2-35M 35M 78.0 0.428
ESM1b-650M 650M 78.9 0.433
ESM2-650M 650M 79.7 0.438
Prot-T5-XL 3B 79.7 0.447
RDKit 20k 76.9 0.421
$T$ 23k 77.1 0.422
ECFP 22k 78.1 0.437
PeptideReactor (Spänig et al., 2021) is a comprehensive benchmark of 50 datasets, aimed at comparing peptide encoding algorithms. To focus solely on the differences in feature engineering, a Random Forest with 100 trees is consistently used as the classifier. No hyperparameter tuning or class weighting is used for the classifier head, only tuning the feature extraction algorithms is allowed. The authors implemented 48 encodings, based on sequence or structure (using 3D structure descriptors). This benchmark allows us to verify how molecular fingerprints compare to a wide variety of previous peptide encodings from bioinformatics.
For this benchmark, we perform hyperparameter tuning on small grids for fingerprints. We make this exception because, despite its name, many datasets in this benchmark consist of quite large proteins. 17 datasets out of 50 have average sequence lengths of more than 50 amino acids (typical upper limit for peptides), and 8 have average lengths of more than 200. Furthermore, other methods also have their parameters tuned here, in order to fairly assess the capability of the encoding method overall, not its default hyperparameters. With 5-fold CV, we tune: ECFP radius (range [2,4]24[2,4][ 2 , 4 ]), Topological Torsion path length (range [4,6]46[4,6][ 4 , 6 ]), RDKit path length (range [7,9]79[7,9][ 7 , 9 ]). They are still very short-range descriptors, but allowed to capture slightly larger substructures.
Since this is the encoding benchmark, we also check the performance of general “FP encoding“, where we treat the fingerprint type as a hyperparameter. This helps us answer the question "How good are encodings based on the topological graph of a molecule?". This is relevant because the two existing feature groups use amino acid sequence (1D encoding) or spatial peptide structure (3D encoding). Furthermore, this is also more realistic, since we select the best method for each dataset. The results are summarized in Table 7. Due to space limitations, here we report only five best structure- and sequence-based encodings, with detailed results in Appendix D.
[…]
Fingerprints are also highly competitive with top sequence-based methods; notably, the ECFP fingerprint alone secures the 4th position in the benchmark, surpassing 45 other encodings. However, we also noticed quite significant differences between fingerprints depending on the dataset. If we consider the general FP encoding, it achieves the SOTA result on this dataset. This means that fingerprints form robust peptide representations that are efficient across a wide variety of properties and protein sizes. We note that the authors of (Spänig et al., 2021) conclude that structural encodings are inferior to sequence-based ones, both in terms of time required and performance. Fingerprint-based encoding constitutes a counterexample to those findings, albeit based on a topological 2D structure, instead of a spatial one.
Table 7: Results on PeptideReactor benchmark (Spänig et al., 2021).
Model group Model Average F1
Structure disord 53.5
electr 60.9
distan 61.3
delaun 69.5
qsar 72.8
Sequence dde 81.3
ngram_ 81.6
psekraac 81.8
dist_f 82.2
cksaap 82.4
Fingerprint RDKit 80.6
$T$ 80.6
ECFP 81.7
FP encoding 82.9
5 Conclusions
We presented an approach to peptide function prediction using count-based variants of hashed molecular fingerprints. They utilize atomic-level representation of peptides using their molecular graphs, which is cheap to obtain. This feature extraction method is also extremely fast to compute. Despite being strictly local, short-range, and generally simple descriptors, they show remarkable performance across a wide range of tasks.
Combined with the LightGBM classifier, this approach achieves superior results in a comprehensive evaluation of six benchmarks consisting of 126 datasets. In addition to classification quality, they are highly performant and parameter-efficient. They outperform a wide variety of other approaches, including GNNs, pretrained amino acid sequence transformers, feature engineering, and multimodal models.
This work also constitutes, to the best of our knowledge, the largest comparison of fingerprints applied to peptides in the literature. In contrast to previous bioinformatics works, we use count variants, instead of binary fingerprints, and show that this detail is crucial to obtain a strong encoding. Our major methodological contribution is showing that these inherently short-range models obtain SOTA results on peptide tasks from Long Range Graph Benchmark (LRGB). ECFP-based classifier with radius 2 outperforms the previous best model by 1.5% AUPRC on Peptides-func. These results challenge the assumed importance of long-range dependencies on this task, further strengthening previous observations from the literature (Tönshoff et al., 2024).
In future work, we plan to extend this approach to property prediction of larger proteins. We also want to analyze fingerprint performance more thoroughly, based on explainable AI approaches, and to provide more insights into the nature of short- and long-range dependencies in peptides.
We conclude that atom-level, short-range approaches based on molecular fingerprints are strong models for peptide property prediction. The use of such models constitutes an important element in verifying claims such as long-range dependencies on molecular tasks.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional benchmark metrics
Here, we report additional metrics for AMPs benchmarks from (Gao et al., 2024) (Table 8) and (Xu et al., 2021) (Table 9), as well as AutoPeptideML benchmark (Fernández-Díaz et al., 2024) (Table 10). They were omitted in the main text due to space limitations.
Table 8: Results for BERT benchmark (Gao et al., 2024) results.
Metric Model ADAPTABLE APD CAMP dbAMP DRAMP YADAMP avg
Recall AMP-BERT 67.60 79.40 80.40 54.20 54.70 82.60 69.82
Bert-Protein 58.40 94.10 76.30 83.80 47.60 47.60 67.97
cAMPs_pred 44.50 58.80 57.90 36.30 24.70 73.90 49.35
LM_pred 40.00 55.90 49.10 45.10 31.80 69.60 48.58
LM_pred (BFD) 64.60 88.20 78.00 58.50 52.80 82.60 59.67
ECFP 75.60 97.10 88.70 95.10 66.80 100.00 87.22
$T$ 75.60 97.10 89.10 94.00 67.10 100.00 87.15
RDKit 77.20 97.10 88.70 94.00 66.50 100.00 87.25
Precision AMP-BERT 80.20 77.10 84.20 74.40 77.10 82.60 79.27
Bert-Protein 79.20 86.50 85.40 82.10 77.10 77.10 81.23
cAMPs_pred 98.20 100.00 97.90 94.50 94.80 100.00 97.57
LM_pred 92.50 100.00 94.00 92.80 94.00 100.00 95.55
LM_pred (BFD) 86.30 91.20 90.20 83.40 83.90 86.40 91.62
ECFP 86.90 89.20 87.70 88.20 89.00 92.00 88.83
$T$ 86.40 91.70 88.50 87.80 89.80 92.00 89.37
RDKit 87.50 89.20 88.00 89.00 89.80 85.20 88.12
AUROC AMP-BERT 81.70 91.80 90.60 78.40 76.90 90.70 85.02
Bert-Protein 79.40 96.10 90.90 91.90 74.00 74.00 84.38
cAMPs_pred 82.30 91.80 87.00 81.90 68.10 93.00 84.02
LM_pred 88.00 97.20 92.30 89.50 82.80 98.70 91.42
LM_pred (BFD) 86.30 97.40 93.60 87.60 83.90 93.40 90.47
ECFP 89.50 98.40 94.30 97.20 87.80 100.00 94.53
$T$ 89.30 98.20 94.30 97.10 87.60 100.00 94.42
RDKit 90.20 99.30 94.80 97.00 88.20 100.00 94.92
F1 AMP-BERT 81.70 78.30 82.30 62.70 64.00 82.60 75.27
Bert-Protein 79.40 90.10 80.60 82.90 58.80 58.80 75.10
cAMPs_pred 61.30 74.10 72.80 52.40 39.20 85.00 64.13
LM_pred 55.90 71.70 64.50 60.70 47.60 82.10 63.75
LM_pred (BFD) 73.90 89.60 83.60 68.70 64.80 84.40 70.32
ECFP 80.80 93.00 88.20 91.50 76.30 95.80 87.60
$T$ 80.60 94.30 88.80 90.80 76.90 95.80 87.87
RDKit 82.00 93.00 88.40 91.40 76.40 92.00 87.20
Table 9: Additional metrics for results on (Xu et al., 2021).
Method Accuracy AUROC F1 MCC Recall Specificity
AMPscannerV2 56.8 58.5 54.8 0.137 52.3 61.3
iAMP-2L 59.2 59.2 36.8 0.261 23.8 94.7
ADAM-SVM 61.2 61.2 47.1 0.264 34.6 87.8
ampir 56.3 61.9 37.9 0.156 26.6 85.9
MLAMP 55.3 62.9 23 0.194 13.3 97.2
ADAM-HMM 68.4 68.4 62.3 0.39 52.1 84.7
AMPlify 64.2 69.7 46.2 0.381 30.7 97.6
AMPEP 65.8 72.7 48.7 0.425 32.5 99.2
AMPfun 67.4 73.5 55.5 0.414 40.6 94.3
RDKit 67.5 73.7 58.5 0.388 45.8 89.1
ECFP 69.3 75.3 60.9 0.426 47.9 90.6
$T$ 69.4 74.9 61.7 0.424 49.3 89.5
Table 10: Additional metrics for results on AutoPeptideML (Fernández-Díaz et al., 2024).
Model # params Accuracy MCC AUROC F1
ProtBERT 420M 68.5 0.375 75.9 67.7
ESM2-150M 150M 69.8 0.402 77.7 68.5
Prost-T5 3B 70.0 0.409 77.1 69.0
ESM2-8M 8M 70.2 0.418 77.5 69.4
ESM2-35M 35M 71.0 0.428 78.0 70.2
ESM1b-650M 650M 71.1 0.433 78.9 69.2
ESM2-650M 650M 68.0 0.438 79.7 69.7
Prot-T5-XL 3B 68.9 0.447 79.7 70.4
RDKit 20k 70.7 0.421 76.9 69.6
$T$ 23k 70.8 0.422 77.1 69.4
ECFP 22k 71.5 0.437 78.1 70.2
Appendix B Additional classifiers on LRGB
Here, we present results of two additional classifiers on the LRGB benchmark datasets: Random Forest (Breiman, 2001) and Extremely Randomized Trees (Geurts et al., 2006). For both we use 500 trees, entropy or squared error as cost function (for classification and regression, respectively), and class weighting for classification. We can report standard deviations for them over 10 random seeds, because they are nondeterministic.
Peptides-func results are in Table 11, and Peptides-struct in Table 12. For clarity, we also include LightGBM results from the main body, which is a deterministic classifier and thus does not have standard deviation.
Random Forest and Extremely Randomized Trees classifiers have very low standard deviation, highlighting the stability of the fingerprint-based approach. Models based on ECFP features consistently achieve the best results in all cases, but other fingerprints and classifiers also obtain results highly competitive with long-range GNNs.
Table 11: AUPRC $\uparrow$ for different classifiers on Peptides-func.
LightGBM Random Forest Extremely Randomized Trees
[…]
Table 12: MAE $\downarrow$ for different classifiers on Peptides-struct.
LightGBM Random Forest Extremely Randomized Trees
[…]
Appendix C Additional LRGB timings
Here, we present expanded timings of fingerprints and compare them to times reported in LRGB (Dwivedi et al., 2022) (Appendix C.2), summarizing them in Table 13. Times for fingerprint-based models are an average of 10 runs on 12-core Intel Core i7-12700KF. We measure them for Peptides-func, and for Peptides-struct results were almost identical. Timings for GNNs are taken from (Dwivedi et al., 2022), and use Nvidia A100 GPU.
The entire time for our approach (feature extraction + classifier training) is shorter than precomputing structural embeddings or training even a single epoch of SAN model. Graph transformer model has faster epochs, they still require precomputing LapPE embeddings, while also giving much worse results than fingerprints (see main body for results table).
One should also take into consideration the difference in raw compute power (CPU vs powerful GPU) required to get those times, which makes the difference even more significant.
Table 13: Time of computation on Peptides-func.
Model Time [s]
Transformer+LapPE 5.9
SAN+LapPE 53.6
SAN+RWSE 49.7
LapPE encoding 74
RWSE encoding 53
ECFP 19
$T$ 15.2
RDKit 42.8
Appendix D Additional PeptideReactor results
Here, we present the performance of all encodings on PeptideReactor (Spänig et al., 2021), in Table 14.
Table 14: Full results on PeptideReactor benchmark (Spänig et al., 2021).
Encoding Avg F1 Type
FP encoding 82.9 fingerprint
cksaap 82.4 sequence
dist_f 82.2 sequence
psekraac 81.8 sequence
ECFP 81.7 fingerprint
ngram_ 81.6 sequence
dde 81.3 sequence
dpc 80.6 sequence
$T$ 80.6 fingerprint
RDKit 80.6 fingerprint
fldpc_ 79.4 sequence
qsorde 78.9 sequence
waac_a 78.8 sequence
apaac_ 78.2 sequence
aac 78.1 sequence
binary 77.9 sequence
ctdd 77.9 sequence
paac_l 77.9 sequence
aainde 77.8 sequence
tpc 76.6 sequence
cksaag 76.5 sequence
ctriad 75.9 sequence
ctdt 75.7 sequence
ctdc 75.6 sequence
ksctri 75.6 sequence
gtpc 75 sequence
gdpc 74 sequence
flgc_a 73.8 sequence
fft_aa 73.3 sequence
qsar 72.8 structure
nmbrot 72.3 sequence
zscale 72.1 sequence
eaac_w 71.6 sequence
blomap 70.9 sequence
egaac_ 70.3 sequence
blosum 69.8 sequence
delaun 69.5 structure
gaac 69.5 sequence
moran_ 69 sequence
socnum 69 sequence
geary_ 68.9 sequence
cgr_re 68.8 sequence
distan 61.3 structure
electr 60.9 structure
disord 53.5 structure
sseb 51.7 structure
asa 51.5 structure
ta 51.4 structure
ssec 48.3 structure